OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (-12,10) and is perpendicular to 4x + 6y = 3

OpenStudy (anonymous):

y=3/2x+28

OpenStudy (anonymous):

3/2 is not perp to 2/3

OpenStudy (anonymous):

wouldn't the slope be positive because the slope of the other line is -2/3. and in order to be perpindicular you have to have the opposite reciprocal slope.

OpenStudy (anonymous):

The slope is -2/3 because its 6y=-4x+3, then y=-2/3x+1/2. This makes the perpendicular 3/2.

OpenStudy (anonymous):

whoops my bad. careless error. i forgot it was -4x. then u are correct.

OpenStudy (anonymous):

3/2 is perp to -2/3